Hello,
it has been a while since I did any linear equations and I would appreciate if you can help me with this questions:
Solve the following system of linear equations to find unique values for A and B?
i) A=1600-400B
ii) A=-200+500B
Thank you so much,
Enna

Answer by rahman(247)

You can modify both of system of linear equations:
i) A=1600-400B ------------- A + 400B = 1600
ii) A=-200+500B ------------ A - 500B = -200
You can solve with elimination method:
A + 400B = 1600
A - 500B = -200 -
----------------
900B = 1800
B = 2
A + 400B = 1600 ------- A + 400(2) = 1600 ----- A = 1600 - 800 = 800
So, we find the solution of system of linear equations: A = 800 & B = 2
Ok